Suppose Praxis Corporations CFO is evaluating a project with the following cash inflows. She does not know the projects initial cost; however, she does know that the projects regular payback period is 2.5 years If the projects weighted average cost of capital (WACC) Year Cash Flow Year 1 $350,000 Year 2 $425,000 Year 3 $425,000 Year 4 $450,000 is 796, what is its NPV? $481,253 $380,992 $401,044 Xⓔ $461,201
Answer #1

Payback period=Last period with a negative cumulative cash flow+(Absolute value of cumulative cash flows at that period/Cash flow after that period).

Hence initial outlay=(350,000+425000+(425000*0.5))=$987500

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=350,000/1.07+425000/1.07^2+425000/1.07^3+450,000/1.07^4

=$1388543.71

NPV=Present value of inflows-Present value of outflows

=$1388543.71-$987500

=$401044(Approx).